A merchant mixed 10 lb of a cinnamon tea with 5 lb of spice tea. The 15-pound mixture cost $40. A second mixture included 12 lb of the cinnamon tea and 8 lb of spice tea. The 20 lb mixture cost $54. Find the cost per pound of the cinnamon tea and the spice tea.
Answer by lwsshak3(3711)

A merchant mixed 10 lb of a cinnamon tea with 5 lb of spice tea. The 15-pound mixture cost $40. A second mixture included 12 lb of the cinnamon tea and 8 lb of spice tea. The 20 lb mixture cost $54. Find the cost per pound of the cinnamon tea and the spice tea.
**
let x=cost per pound of cinnamon
let y=cost per pound of spice tea
..
10x+5y=40
12x+8y=54
..
80x+40y=320 (mult. by 8)
60x+40y=270 (mult. by 5)
subtract
20x=50
x=50/20=2.5
..
5y=40-10x=40-25=15
y=3
ans:
cost per pound of cinnamon=$2.50
cost per pound of spice tea=$3.00